# ADVANCING MATHEMATICS RESEARCH WITH LARGE LANGUAGE MODELS

### v1

@@ Abstract.
The main drawback of using generative AI for advanced mathematics via Large Language Models (LLMs) is that they are probabilistic pattern-matchers, not logical reasoning engines. However, LLMs can pick up on patterns in higher mathematics that are difficult for humans to see. By putting the design of LLMs to their advantage, mathematicians may use them as powerful interactive assistants that can carry out laborious tasks, generate and debug code, check examples, formulate conjectures and more. We discuss how LLMs can be used to advance mathematics research by careful use of prompt engineering. We also discuss the integration of LLMs with Computer Algebra Systems and formal proof assistants such as Lean.
@@ 1. Introduction
Mathematicians have mixed views about the role of LLMs in the mathematical landscape. While these models can look convincingly like they replicate known mathematics, on careful scrutiny, they fall short of the standards of rigor that the field requires.
@@
In Section 2, we discuss the mechanics of probabilistic language generation. In Sections  37, we explore the foundational capabilities and limitations of LLMs, the nature of their training data, and their ability to detect geometric patterns in mathematical language. In Sections 815, we discuss applications and user techniques. This includes examples from combinatorial group theory, a comparison of current models, and a guide to prompt engineering, fact-checking and managing the context window. In Section 16 we discuss the integration of LLMs with Computer Algebra Systems and formal proof assistants like Lean. In Section 17, we give some examples of advancements in research in infinite dimensional group theory obtained with the strategic use of hybrid methods.
@@ 2.1. The LLM as a probabilistic knowledge graph
Consider the following simplified view of an LLM as a ‘knowledge graph’ which represents all possible connections and probabilities within the LLM’s knowledge base.
@@
An LLM generates text by predicting the next word, while calculating the probabilities of all possible next nodes. It chooses one of the most likely paths through the knowledge graph.
@@ 2.2. Tokens: the building blocks of language
Let $V$ be a finite vocabulary consisting of token representations of the training data for an LLM. An input to an LLM is a sequence of tokens $(t_{1},t_{2},\dots,t_{n})$. An output of an LLM is a sequence of tokens $(t_{1},t_{2},\dots,t_{m})$, where each $t_{i}\in V$.
@@
All AI language models have a maximum token limit for input and output.
Model Maximum Input Limit (tokens) Maximum Output Limit (tokens)
Gemini 1,048,576 65,536
ChatGPT 272,000 128,000
Claude 200,000 32,000
The token input limit has an impact on the work possible in a single chat session.
@@ 3. Drawbacks of using generative AI in mathematics
The primary drawback of using current LLMs for mathematics is that they are probabilistic pattern-matchers, not logical reasoning engines. Their core function is to predict the next most likely word or symbol in a sequence, not to apply deterministic mathematical rules. An LLM is an expert at reconstructing what a proof should ‘look like’, based on every proof it has ever seen. This limitation leads to several key problems:
Mathematical hallucinations: LLMs generate outputs that look plausible and are often formatted correctly but contain nonsensical logic, invented theorems, or critical errors. They can state these falsehoods with the same confident authority as factual information.
@@
Propagation of errors: LLMs learn from their training data, which includes errors and misconceptions found online. For example, it is known that ChatGPT 4.0 contains training data from retracted scientific papers [MITTR2025]. An LLM will inherit and reproduce these mistakes, making them unreliable for tasks that require rigor.
@@
4. Mathematical training data for LLMs
Training data for LLMs includes web documents, code with mathematical content, textbooks, online faculty-authored lecture notes and course materials, solutions to problem sets, as well as academic and scientific journal papers and content from arXiv. Google’s data collection from Google Books, for example, is unparalleled in its scale.
@@ 4. Mathematical training data for LLMs
A key limitation is that an LLM’s knowledge base is heavily biased towards materials that are easily digitized.
5. LLMs simulate computer algebra systems
LLMs simulate a scientific computing environment by delegating the computation to specialized Python libraries.
Table 1. Key Python libraries used by LLMs for mathematics
@@ 5. LLMs simulate computer algebra systems
NumPy Numerical Computing Handles arrays & matrices. Used by LLMs for linear algebra (matrix multiplication, row reduction, eigenvalues).
@@
6. How much do LLMs ‘know’?
Current LLMs are trained on human data, which means they cannot generate knowledge outside of that data. While they can go beyond the knowledge of any single human, they can only come up with new ideas through extrapolation, not discovery from first principles. Their core capability is Natural Language Reasoning - the ability of LLMs to process, synthesize and generate human language by identifying and replicating statistical patterns from vast amounts of data in the form of tokens. In contrast, Symbolic Deduction is manipulation based on formal rules applied to abstract symbols, as in the operation of a Computer Algebra System. Current LLMs are built to imitate reasoning based on patterns, but are incapable of performing true symbolic deduction.
However, just as mathematicians are becoming familiar with the use of LLMs, their internal architecture is already changing. In 2025, Google’s DeepMind achieved a gold-medal level score alongside the International Mathematical Olympiad using their AlphaProof and AlphaGeometry systems. These systems did not use a separate formal proof assistant. The key innovation in their new model is the integration of Symbolic Deduction with Natural Language Reasoning, representing a first step towards an AI-based alternative to formal proof assistants. Their new models also incorporate techniques like parallel thinking, which allowed them to explore multiple solutions simultaneously.
The significant advancement in reasoning capability has been achieved by an advanced form of chain-of-thought and tree-of-thought reasoning. These are important tools in when using LLMs for mathematics. We give some examples in Section 12.
@@ 6. How much do LLMs ‘know’?
7. LLMs and patterns in higher mathematics
An LLM can identify mathematical patterns that are imperceptible to humans. It performs massive scale statistical analysis on the symbolic structure of mathematics itself. It can find cross-disciplinary correlations and potential analogies that humans may not think of. It is unconstrained by human cognitive biases and the boundaries of human knowledge.
The way mathematicians are trained to think assumes the progression of ideas of mathematical discovery. We have learned concepts in a chronological sequence. LLMs do not ‘know’ that mathematical fields developed sequentially. For an LLM, mathematical concepts coexists simultaneously. They are linked only by the statistical and structural patterns embedded in the language of mathematics. Due to their inherent design, they can propose unexpected pathways, connections and points of view.
@@ 7.1. Advantages of this capability:
An LLM sees mathematics as a language with a complex grammar and vocabulary built up of symbols and diagrams. It learns the statistical relationships between the tokens in this language.
Trained on the entire available library of mathematics and science without disciplinary boundaries, it learns the statistical relationships between tokens, allowing it to explore the entire space of mathematical connections, not just those that individual mathematicians are trained to look for.
LLMs don’t ‘see’ symbols and equations. There is a map $\varphi:V\to\mathbb{R}^{d}$ (for $d\leq 13,000$) from the vocabulary of tokens to a high-dimensional vector space. The dimension $d\sim|V|^{1/4}$ is a model parameter. Relationships in the vocabulary are translated into geometric relationships in $\mathbb{R}^{d}$.
The model can detect geometric relationships in $\mathbb{R}^{d}$ that are impossible for humans to visualize and that are not obvious from the symbolic definitions. For example, it uses a sophisticated method of projecting tokens onto subspaces via linear transformations to isolate a specific feature to look at. It then reviews all isolated facts to decide on an output.
It is this ability to perceive and analyze non-obvious geometric patterns that allows LLMs to generate novel conjectures and insights when prompted appropriately.
@@ 7.2. Human-LLM collaborative research
Using an LLM as a collaborative research partner, rather than as a search engine or text generator, involves a continuously refined, iterative dialogue. The researcher guides the model’s output through a series of increasingly specific constraints, corrections, and questions, often in a loop that involves downloading, correcting, and re-uploading the model’s work for further prompting.
For open research questions, this can be a lengthy process, sometimes involving hundreds of prompts in a single chat, to obtain a final product that is useful. A key technique is ‘in-context learning’, where researchers provide background information like papers or books via prompts. The LLM keeps this material in its short-term memory, becoming more adept at handling the research topic as the conversation history forms an expanding context window.
@@ 8. Why AI is a useful tool in combinatorial group theory
Computer algebra systems such as GAP, Magma, and Sage can solve the word problem for specific families of groups using coset rewriting methods and structural properties. For new groups whose structure is not well understood, an LLM can be used as an interactive research assistant.
A simple example: let $G=BS(1,2)=\langle x,y\mid x^{-1}yx=y^{2}\rangle$. This is the so-called Baumslag-Solitar group $BS(1,2)$. It’s word problem is known to be solvable.
@@
Solve the word problem for the word $w=yxy^{-2}x^{-1}$:
@@
So the word $w=yxy^{-2}x^{-1}$ represents the identity element in $G$.
@@ 8.1. Solving algorithmic research questions
However, this is not true symbolic manipulation: LLMs mimic the process through pattern recognition, so their output must always be verified and corrected.
In addition, LLMs can generate code for computer algebra systems such as GAP, Magma, and SageMath with varying degrees of reliability. Proficiency is particularly high with SageMath, given its Python-based syntax and integration with GAP for computations in group theory and other areas.
9. Examples of the use of LLMs to address pattern-matching problems in higher mathematics
You can ask an LLM to:
@@ 10. Comparison of LLMs
For the most difficult mathematical tasks, Gemini, ChatGPT, and Claude are all roughly in the same class with respect to performance.
Though Gemini models perform slightly better in advanced mathematical benchmarks, as of 2024.
@@
‘Advanced mathematics’ refers to the standard benchmarks for competition-level mathematics, testing symbolic reasoning in algebra, calculus, number theory, etc.
Advantages of ChatGPT are its plug-ins: WolframAlpha can be enabled in the professional version. Another plug-in called Sci-space has access to 300 million peer-reviewed journal papers.
@@ 11. General rules for writing prompts for advanced math questions:
You can also ask an LLM to generate a prompt for a given task.
12. Prompts for advanced mathematics
12.1. Chain-of-thought prompt
This refers to guiding the LLM to think step-by-step, generating intermediate reasoning before arriving at the final answer [IBMToT]. This type of structured prompt imposes a logical framework on the LLM’s probabilistic generation process. If the LLM shows its work, this reduces errors.
@@ 12.2. Tree-of-thought prompt
This is guiding an LLM to explore multiple different deduction paths simultaneously such that it evaluates its own progress at each step and pursues the most promising path [IBMToT].
@@
These types of example prompts have been shown to be useful in guiding an LLM how to approach complex mathematical questions.
@@ 13. Fact-checking: retrieval-augmented generation
The default output of an LLM is an output that is based on its training data. The ability to access the internet to fact-check in real time is a feature built on top of an LLM, through a mechanism called ‘retrieval-augmented generation’. It may require a specific prompt in order to invoke this feature. Users can also explicitly request a list of websites and references used.
@@
14. A single chat is a ‘context window’
An LLM’s short-term memory is its context window, which consists of a single chat session. This window contains the entire conversation history, including uploaded documents and any corrections made by the user. Its capacity is measured in tokens: while a 120K window can hold a novel, models like Gemini offer an unprecedented 1 million tokens.
However, this memory is finite. Once the token limit is reached, the LLM will begin to forget the earliest messages in the conversation. Work on a difficult mathematics problem must be contained within a single chat: the LLM needs all information and previous steps to be present simultaneously to solve a complex problem.
Since an LLM has no inherent memory between separate sessions, new chats must be initiated with any updated information and results.
15. Other ways to influence the output from an LLM

### v2

# ADVANCING MATHEMATICS RESEARCH WITH GENERATIVE AI

@@ Abstract.
The main drawback of using generative AI models for advanced mathematics is that these models are not logical reasoning engines. However, Large Language Models, and their refinements, can pick up on patterns in higher mathematics that are difficult for humans to see. By putting the design of generative AI models to their advantage, mathematicians may use them as powerful interactive assistants that can carry out laborious tasks, generate and debug code, check examples, formulate conjectures and more. We discuss how generative AI models can be used to advance mathematics research. We also discuss their integration with Computer Algebra Systems and formal proof assistants such as Lean.
@@ 1. Introduction
Mathematicians have mixed views about the role of generative AI models in the mathematical landscape. While these models, such as Large Language Models (LLMs), can look convincingly like they replicate known mathematics, on careful scrutiny, it is apparent that they are just masters of the rhetoric of mathematics. They don’t meet the standards of rigor that the field requires.
@@
The AI landscape is rapidly evolving beyond standard LLMs, with Large Reasoning Models (LRMs) and Large Context Models (LCMs) emerging as the next generation of this technology. While built on the foundation of LLMs, these models have distinct architectures and methods of training (see Section 4).
In Section 2, we discuss the mechanics of probabilistic language generation. In Section 3, we discuss how to manage the context window. In Sections  59, we describe the capabilities and limitations of LLMs, the nature of their training data, and their ability to detect geometric patterns in high-dimensional representations of mathematical language. In Sections 1016, we discuss applications and user techniques. This includes examples from combinatorial group theory, a comparison of current models, a guide to prompt engineering, and fact-checking. In Section 17 we discuss the integration of LLMs with Computer Algebra Systems (CAS) and formal proof assistants like Lean.
@@ 2.1. The LLM as a probabilistic knowledge graph
Consider the following simplified view of a generative AI model as a ‘knowledge graph’ which represents all possible connections and probabilities within the LLM’s knowledge base.
@@
An LLM generates text by predicting the next word, starting from the prompt, while calculating the probabilities of all possible next nodes. It then chooses one of the most likely paths through the knowledge graph.
@@ 2.2. Tokens: the building blocks of language
Let $V$ be a finite vocabulary consisting of token representations of the training data for an LLM. An input to an LLM is a sequence of tokens $(t_{1},t_{2},\dots,t_{n})$. An output is a sequence of tokens $(t_{1},t_{2},\dots,t_{m})$, where each $t_{i}\in V$.
@@
All AI language models have a maximum token limit for input and output. The token input limit has an impact on the work possible in a single chat session.
@@ 2.3. How the prompt shapes an LLM’s output
3. A single chat is a ‘context window’
A generative AI model’s short-term memory is its context window, which consists of a single chat session. This window contains the entire conversation history, including uploaded documents and any corrections made by the user. Its capacity is measured in tokens: while a 120K window can hold a novel, models like Gemini 2.5 Pro offer an unprecedented 1 million tokens.
However, this memory is finite. Once the token limit is reached, the AI model will begin to forget the earliest messages in the conversation. Work on a difficult mathematics problem must be contained within a single chat: the AI model needs all information and previous steps to be present simultaneously to solve a complex problem.
Since a generative AI model has no inherent memory between separate sessions, new chats must be initiated with any updated information and results.
4. Large Reasoning Models and Large Context Models
Large Reasoning Models (LRMs), also known as a Reasoning Language Model (RLMs), are a class of LLMs trained to carry out tasks that require multi-step deduction, critical thinking, and structured problem-solving. LRMs ‘think’ before they respond.
They deliberate, explore various solution paths, evaluate intermediate steps and revise their reasoning.
LRMs integrate frameworks that support reasoning structures, such as ‘chain-of-thought’ and ‘tree-of-thought’ structures into their systems. We give some examples in Section 14.
The training of LRMs diverges from that of LLMs. They utilize a technique called ‘process supervision’. Unlike the training process of LLMs that only rewards a correct final answer, process supervision rewards the correctness of the intermediate reasoning steps.
Large Context Models (LCMs) are defined by their large context windows and the volume of information that they can handle. This allows them to accept and analyze large datasets, entire codebases, or multimodal inputs, without requiring external retrieval systems. Their training data includes long-form content to improve their ability to track long-range dependencies between input tokens.
These advanced capabilities are becoming integrated into leading AI platforms. The cost, however, is additional computing power and increased time for responses.
Google’s Gemini models, particularly the 2.5 series, demonstrate both LRM and LCM characteristics. The models also use internal ‘thinking processes’ for improved reasoning.
ChatGPT has a ‘reasoning mode’ in its model OpenAI o3. Reinforcement learning was used to teach OpenAI o3 to ‘think’ before generating answers, using what OpenAI refers to as ‘private chain-of-thought’ processes built into their system prompts. This allows the model to plan ahead and reason through tasks, performing a series of intermediate reasoning steps to assist in solving the problem.
Claude models use ‘extended thinking’ or the ability to ‘think out loud’. As LCMs, they offer substantial context windows
@@ 5. Drawbacks of using generative AI in mathematics
The primary drawback of using current AI models for mathematics is that they are probabilistic pattern-matchers, not logical reasoning engines. Their core function is to predict the next most likely word or symbol in a sequence, not to apply deterministic mathematical rules. Generative AI models are experts at reconstructing what a proof should ‘look like’. This limitation leads to several key problems:
Mathematical hallucinations: AI models generate outputs that look plausible and are often formatted correctly but contain nonsensical logic, invented theorems, or critical errors. They can state these falsehoods with the same confident authority as factual information.
@@
Propagation of errors: AI models learn from their training data, which includes errors and misconceptions found online. For example, it is known that ChatGPT 4.0 contains training data from retracted scientific papers [MITTR2025]. An LLM will inherit and reproduce these mistakes, making them unreliable for tasks that require rigor.
@@
6. Mathematical training data for generative AI models
Training data for generative AI models includes web documents, code with mathematical content, textbooks, online faculty-authored lecture notes and course materials, solutions to problem sets, as well as academic and scientific journal papers and content from arXiv. Google’s data collection from Google Books, for example, is unparalleled in its scale.
@@ 6. Mathematical training data for generative AI models
A key limitation is that a generative AI model’s knowledge base is heavily biased towards materials that are easily digitized.
7. Generative AI models simulate computer algebra systems
Advanced generative AI models simulate a scientific computing environment by delegating the computation to specialized Python libraries.
Table 1. Key Python libraries used by generative AI models for mathematics
@@ 7. Generative AI models simulate computer algebra systems
NumPy Numerical Computing Handles arrays & matrices. Used by generative AI models for linear algebra (matrix multiplication, row reduction, eigenvalues).
@@
8. How much do generative AI models ‘know’?
Current generative AI models are trained on human data, which means they cannot generate knowledge outside of that data. While they can go beyond the knowledge of any single human, they can only come up with new ideas through extrapolation, not discovery from first principles. Their core capability is Natural Language Reasoning - the ability of generative AI models to process, synthesize and generate human language by identifying and replicating statistical patterns from vast amounts of data in the form of tokens.
In contrast, Symbolic Deduction is manipulation based on formal rules applied to abstract symbols, as in the operation of a Computer Algebra System. Current generative AI models are built to imitate reasoning based on patterns, but are incapable of performing true symbolic deduction.
However, just as mathematicians are becoming familiar with the use of generative AI models, their internal architecture is already changing.
In 2024, Google’s DeepMind reported silver-medal level on International Mathematical Olympiad problems using AlphaProof and AlphaGeometry. In 2025, DeepMind reported gold-medal level using Gemini ‘Deep Think’. These systems did not use a separate formal proof assistant. The key innovation in their new model is the integration of Symbolic Deduction with Natural Language Reasoning, representing a first step towards an AI-based alternative to formal proof assistants. Their new models also incorporate techniques like parallel thinking, which allowed them to explore multiple solutions simultaneously.
The significant advancement in reasoning capability has been achieved by an advanced form of chain-of-thought and tree-of-thought reasoning. These are important tools when using generative AI models for mathematics. We give some examples in Section 14.
@@ 8. How much do generative AI models ‘know’?
9. generative AI models and patterns in higher mathematics
A generative AI model can identify mathematical patterns that are imperceptible to humans. It performs massive scale statistical analysis on the symbolic structure of mathematics itself. This allows it to simulate a mathematician’s intuition. It can find cross-disciplinary correlations and potential analogies that humans may not think of. It is unconstrained by human cognitive biases and the boundaries of human knowledge.
The way mathematicians are trained to think assumes the progression of ideas of mathematical discovery. We have learned concepts in a chronological sequence. AI models do not ‘know’ that mathematical fields developed sequentially. For an AI model, mathematical concepts coexist simultaneously. They are linked only by the statistical and structural patterns embedded in the language of mathematics. Due to their inherent design, they can propose unexpected pathways, connections and points of view.
@@ 9.1. Advantages of this capability:
A generative AI model sees mathematics as a language with a complex grammar and vocabulary composed of symbols and diagrams. It learns the statistical relationships between the tokens in this language.
Trained on the entire available library of mathematics and science without disciplinary boundaries, an LLM learns the statistical relationships between tokens, allowing it to explore the entire space of mathematical connections, not just those that individual mathematicians are trained to look for.
Generative AI models don’t ‘see’ symbols and equations. There is a map $\varphi:V\to\mathbb{R}^{d}$ (for $d\leq 13,000$) from the vocabulary $V$ of tokens to a high-dimensional vector space (the embedding space). The dimension $d\sim|V|^{1/4}$ is a model parameter. Relationships in the vocabulary are translated into geometric relationships in $\mathbb{R}^{d}$ [mikolov2013].
Vocabulary VVmanwomanqueenkingEmbedding Space $\mathbb{R}^{d}$$\varphi(\text{man})$$\varphi(\text{king})$$\varphi(\text{woman})$$\varphi(\text{queen})$$\vec{v}_{\text{royal}}$$\vec{v}_{\text{royal}}$$\vec{v}_{\text{gender}}$$\vec{v}_{\text{gender}}$$\varphi:V\to\mathbb{R}^{d}$
The model can detect geometric relationships in $\mathbb{R}^{d}$ that are impossible for humans to visualize and that are not obvious from the symbolic definitions.
The image $\varphi(w)$ of a word $w$ doesn’t have a single location in the embedding space. Its ‘meaning’ is its relationship to all other images of tokens in the space and their locations. Furthermore, this meaning changes with context. The model applies a linear transformation $T$ to $\varphi(w)\in\mathbb{R}^{d}$ to project it onto a subspace relevant to its contextual role. This allows it to isolate a specific feature to look at. It then reviews all isolated facts and uses piecewise linear maps to put these facts together and compile an output.
This ability to perceive and analyze non-obvious geometric patterns allows the AI model to generate novel conjectures and insights when suitably prompted.
@@ 9.2. Human-LLM collaborative research
Using a generative AI model as a collaborative research partner, rather than as a search engine or text generator, involves an iterative dialogue that must be repeatedly corrected and refined. The user can improve the model’s output through a series of increasingly specific constraints, corrections, and questions, often in a loop that involves downloading, correcting, and re-uploading the model’s work for further prompting.
For open research questions, this can be a lengthy process, sometimes involving hundreds of prompts in a single chat, to obtain a final product that is useful. A key technique is ‘in-context learning’, where researchers provide background information like papers or books via prompts. The LLM keeps this material in its short-term memory. It becomes more adept at handling the research topic as the conversation history forms an expanding context window.
@@ 10. Why AI is a useful tool in combinatorial group theory
Computer Algebra Systems such as GAP, Magma, and SageMath can solve the word problem for specific families of groups using rewriting methods and structural properties.
A simple example: let $G=BS(1,2)=\langle x,y\mid x^{-1}yx=y^{2}\rangle$. This is the so-called Baumslag-Solitar group $BS(1,2)$. Its word problem is known to be solvable.
@@
As an example, we solve the word problem for the word $w=yxy^{-2}x^{-1}$:
@@
Thus the word $w=yxy^{-2}x^{-1}$ represents the identity element in $G$.
@@ 10.1. Solving algorithmic research questions
However, this is not true symbolic manipulation: generative AI models mimic the process through pattern recognition, so their output must always be verified and corrected.
In addition, generative AI models can generate code for Computer Algebra Systems such as GAP, Magma, and SageMath with varying degrees of reliability. Proficiency is particularly high with SageMath, given its Python-based syntax and integration with GAP for computations in group theory and other areas.
11. Examples of the use of generative AI models to address pattern-matching problems in higher mathematics
You can ask a generative AI model to:
@@ 12. Comparison of generative AI models
For the most difficult mathematical tasks, Gemini, ChatGPT, and Claude are all roughly in the same class with respect to performance. Though Gemini models performed slightly better in the standard benchmarks for competition-level mathematics, testing symbolic reasoning in algebra, calculus and number theory in 2024.
@@
ChatGPT integrates features like WolframAlpha which can be enabled in the professional version. Another GPT called SciSpace has access to hundreds of millions of peer-reviewed journal papers.
@@ 13. General rules for writing prompts for advanced math questions:
You can also ask a generative AI model to generate a prompt for a given task.
14. Structured prompts and system instructions
Structured prompts impose a logical framework on an LLM’s probabilistic generation process. The following examples of prompts have been shown to be useful in guiding generative AI models on how to approach complex mathematical questions [kojima2022]. These methods are also used internally in advanced LRM models like Gemini and ChatGPT.
14.1. Chain-of-thought reasoning
This refers to guiding the LLM to think step-by-step, generating intermediate reasoning before arriving at the final answer [IBMToT]. If the LLM shows its work, this reduces errors.
@@ 14.2. Tree-of-thought prompt
This refers to guiding a generative AI model to explore multiple different deduction paths simultaneously such that it evaluates its own progress at each step and pursues the most promising path [IBMToT].
@@
The emergence of LRMs and LCMs (Section 4) has altered the practice of prompt engineering.
For LRMs, the need for prompt engineering techniques designed to force reasoning, such as explicit chain-of-thought prompting, is reduced. The reasoning process is handled internally. Simpler, goal-oriented prompts that clearly define the problem may be more effective.
With LCMs, the focus is shifting to ‘context engineering’. The challenge becomes optimizing the content and structure of the information provided in the prompt. This maximizes performance and prevents recall accuracy from decreasing as the context window fills.
@@ 15. Fact-checking: retrieval-augmented generation
The default output of a generative AI model is an output that is based on its training data. The ability to access the internet to fact-check in real time is a feature built on top of an LLM, through a mechanism called ‘retrieval-augmented generation’, or RAG. It may require a specific prompt in order to invoke this feature. Users can also explicitly request a list of websites and references used.
@@
16. Other ways to influence the output from a generative AI model